# FOZO: Forward-Only Zeroth-Order Prompt Optimization for Test-Time Adaptation

### 13 Mar 2026 v2

@@ 2.1 Test-Time Adaptation
Despite significant progress, their fundamental reliance on backpropagation limits deployment in resource-constrained edge devices or black-box models lacking gradient support. This fundamental limitation has driven the development of Forward-only TTA methods based entirely on forward propagation, which has garnered attention due to its high computational efficiency and ease of deployment, typically avoiding iterative optimization of core model parameters. Early works like AdaBN [23] and PredBN [26] primarily focused on BN layer calibration by replacing statistics. More sophisticated strategies include T3A [17], which adapts by adjusting only the classifier head via online-updated ‘pseudo-prototypes’, and LAME [2], which directly corrects output logits online through Laplace Approximated Maximum-a-Posteriori estimation without modifying model parameters. Recently, FOA [30] proposed a novel paradigm for optimizing model parameters without backpropagation, by employing an evolutionary algorithm [13] to update learnable visual prompts at the ViT input layer.
@@
6:   Sample random seed $s$ and generate $\mathbf{Z}\sim\mathcal{N}(0,I_{d})$
@@ 3.1 Prompt Optimization
Visual-Prompt Tuning [18] has demonstrated superior parameter and data efficiency in adapting pre-trained Vision Transformer (ViT) [9] to downstream application domains. We thus aim to leverage its efficiency and develop test-time adaptation method. Specifically, for an NN-layer pre-trained ViT, the input embedding to the first layer, $\mathbf{E}_{0}=\{\mathbf{e}_{0}^{j},j\in\mathbb{N},0\leq j\leq m\}$, comprises a learnable [CLS] token $\mathbf{e}^{0}_{0}\in\mathbb{R}^{d}$ and $m$ embeddings corresponding to image patches. In this setup, a set of learnable $d$-dimensional prompts $\mathbf{P}=\{\mathbf{p}^{k}\in\mathbb{R}^{d}|k\in\mathbb{N},1\leq k\leq p\}$ is inserted into $\mathbf{E}_{0}$ before the first layer $L_{1}$. The subsequent forward pass through the frozen layers is defined as:
@@ 4.2.2 Convergence Rate
We first establish FOZO’ expected descent with a single optimization step and then prove its convergence over iterations.
@@
For a detailed proof of Theorem 1 and 2, please refer to Supplementary B
@@ 4.4.1 Aligning Deep-Shallow Feature Statistics
To achieve robust adaptation, we propose a group-wise Deep-Shallow alignment strategy. This involves collecting [CLS] token activations from both shallow ($e_{1}^{0},\ldots,e_{N/2}^{0}$) and deep ($e_{N/2+1}^{0},\ldots,e_{N}^{0}$) layers of the model during inference and computing their respective means and variances ($\mu^{T}_{k},\sigma^{T}_{k}$) and aligning them with pre-computed source domain statistics ($\mu^{S}_{k},\sigma^{S}_{k}$). Thus the aligning objective $\mathcal{L}_{stats}$ can be expressed as:
$\mathcal{L}_{stats}=\sum_{k\in\{shallow,deep\}}\left(\|{\mu}^{T}_{k}-\mu^{S}_{k}\|_{2}+\|{\sigma}^{T}_{k}-\sigma^{S}_{k}\|_{2}\right)$ (13)
@@ 4.4.2 Entropy Minimization Loss
$\mathcal{L}_{ent}=-\sum_{i=1}^{B}\sum_{a=1}^{K}p_{i,a}\log(p_{i,a})$ (14)
where $B$ is the batch size, $K$ is the number of classes, and $p_{i,a}$ is the predicted probability for class $a$ of sample ii.
@@
$\mathcal{L}=\lambda\mathcal{L}_{stats}+\mathcal{L}_{ent}$ (15)
@@ 5.1.2 Implementation Details
For fair comparison, we set the batch size of test samples to 64. We use 3 prompt embeddings, initialized uniformly. We employ an n-SPSA zero-order gradient estimator, where $n$ is set to 1 as specified in Eqn. 6. The initial learning rate $\eta$ in Eqn. 7 is 0.08. In Eqn. 8, The decay factor $\alpha$ is set to 0.9 , the threshold factor $\tau$ to 1.05, and the moving average factor $\beta$ for the historical average loss to 0.9. The trade-off parameter $\lambda$ in Eqn. 15 is set to 0.4. Following FOA, source training statistics are estimated using the validation set of ImageNet-1K. Please refer to Supplementary A for more implementation details.
@@ 5.4.3 Hyperparameters Analysis
To evaluate FOZO’s sensitivity to key hyperparameters, we conduct more ablation studies. Fig E.1(a) shows minor accuracy variations across 1 to 10 prompts. While 8 prompts yield the highest accuracy (62.5%), practical test-time adaptation often precludes fine-tuning prompts on unseen data. Therefore, we select 3 prompts for our main experiments, balancing performance with practical considerations. Fig E.1(b) reveals the effect of batch size. Smaller batches (e.g., 4, 8, 16) show reduced accuracy (e.g., 60.9% at batch size 4) due to higher variance in statistics and noisy zeroth-order gradient estimates. Performance peaks at batch size 32 (62.3%), benefiting from more adaptation steps. However, increasing to 128 leads to a slight decrease (61.2%), as large batches may over-smooth feature statistics, losing fine-grained domain information. For our main experiments, we select a batch size of 64 to ensure a fair comparison.
6 Conculsion
We propose Forward-Only Zeroth-Order Optimization (FOZO), a novel forward-only TTA paradigm. Compared to the latest forward-only methods, FOZO offers significant advantages: it consistently achieves higher accuracy with notably lower runtime, memory, and number of parameters needing updates. Furthermore, both theoretical analysis and experiments demonstrate that our dynamic perturbation strategy enables faster convergence during adaptation. These advantages demonstrate that FOZO serves as a highly efficient solution, particularly suited for deployment on resource-constrained edge devices and quantized models.

### 26 Mar 2026 v3

@@ 2.1 Test-Time Adaptation
Despite significant progress, their fundamental reliance on backpropagation limits deployment in resource-constrained edge devices or black-box models lacking gradient support. This fundamental limitation has driven the development of Forward-only TTA methods based entirely on forward propagation, which has garnered attention due to its high computational efficiency and ease of deployment, typically avoiding iterative optimization of core model parameters. Early works like AdaBN [23] and PredBN [26] primarily focused on BN layer calibration by replacing statistics. More sophisticated strategies include T3A [17], which adapts by adjusting only the classifier head via online-updated ‘pseudo-prototypes’, and LAME [2], which directly corrects output logits online through Laplace Approximated Maximum-a-Posteriori estimation without modifying model parameters. Recently, FOA [30] proposed a paradigm for optimizing model parameters without backpropagation, by employing an evolutionary algorithm [13] to update learnable visual prompts at the ViT input layer.
@@
6:   Generate $\mathbf{Z}\sim\mathcal{N}(0,I_{d})$ using random seed $s$
@@ 3.1 Prompt Optimization
Visual-Prompt Tuning [18] has demonstrated superior parameter and data efficiency in adapting pre-trained Vision Transformer (ViT) [9] to downstream application domains. We thus aim to leverage its efficiency and develop test-time adaptation method. Specifically, for an $N$-layer pre-trained ViT, the input embedding to the first layer, $\mathbf{E}_{0}=\{\mathbf{e}_{0}^{j},j\in\mathbb{N},0\leq j\leq m\}$, comprises a learnable [CLS] token $\mathbf{e}^{0}_{0}\in\mathbb{R}^{d}$ and $m$ embeddings corresponding to image patches. In this setup, a set of learnable $d$-dimensional prompts $\mathbf{P}=\{\mathbf{p}^{k}\in\mathbb{R}^{d}|k\in\mathbb{N},1\leq k\leq p\}$ is inserted into $\mathbf{E}_{0}$ before the first layer $L_{1}$. The subsequent forward pass through the frozen layers is defined as:
@@ 4.2.2 Convergence Rate
We first establish FOZO’ expected descent with a single step and then prove its convergence over iterations.
@@
For a detailed proof of Theorem 1 and 2, please refer to Supplementary B.
@@ 4.3 Dynamic Perturbation
While Theorem 2 assumes a monotonic decay $\epsilon_{t}\to 0$ for strict asymptotic precision, the dynamic reset in Eq. 12 acts as a domain-aware re-initialization. Under non-stationary TTA streams, FOZO essentially operates as a cascade of convergence processes. Theorem 2 guarantees the performance lower-bound and convergence within each piece-wise stationary domain. Since practical TTA prioritizes rapid adaptation over asymptotic precision, this reset mechanism redirects the model for broader exploration during extreme shifts without sacrificing the optimization progress accumulated in earlier iterations.
@@ 4.4.1 Aligning Deep-Shallow Feature Statistics
To achieve robust adaptation, we propose a group-wise Deep-Shallow alignment strategy. This involves collecting [CLS] token activations from both shallow ($e_{1}^{0},\ldots,e_{N/2}^{0}$) and deep ($e_{N/2+1}^{0},\ldots,e_{N}^{0}$) layers of the model during inference and computing their respective means and variances ($\mu^{T}_{k},\sigma^{T}_{k}$) for each group $k\in\{\mathrm{shallow},\mathrm{deep}\}$, and aligning them with pre-computed source domain statistics ($\mu^{S}_{k},\sigma^{S}_{k}$). Thus the aligning objective $\mathcal{L}_{stats}$ can be expressed as:
$\mathcal{L}_{stats}=\sum_{k}\left(\|{\mu}^{T}_{k}-\mu^{S}_{k}\|_{2}+\|{\sigma}^{T}_{k}-\sigma^{S}_{k}\|_{2}\right).$ (13)
@@ 4.4.2 Entropy Minimization Loss
$\mathcal{L}_{ent}=-\sum_{i=1}^{B}\sum_{a=1}^{K}p_{i,a}\log(p_{i,a}),$ (14)
where $B$ is the batch size, $K$ is the number of classes, and $p_{i,a}$ is the predicted probability for class $a$ of sample $i$.
@@
$\mathcal{L}=\lambda\mathcal{L}_{stats}+\mathcal{L}_{ent},$ (15)
@@ 5.1.2 Implementation Details
For fair comparison, we set the batch size of test samples to 64. We use 3 prompt embeddings, initialized uniformly. The initial learning rate $\eta$ in Eqn. 7 is 0.08. In Eqn. 12, The decay factor $\alpha$ is set to 0.9 , the threshold factor $\tau$ to 1.05, and the moving average factor $\beta$ for the historical average loss to 0.9. The trade-off parameter $\lambda$ in Eqn. 15 is set to 0.4. Following FOA, source training statistics are estimated using the validation set of ImageNet-1K. Please refer to Supplementary A for more implementation details.
@@ 5.4.3 Hyperparameters Analysis
To evaluate FOZO’s sensitivity to key hyperparameters, we conduct more ablation studies. Fig E.1(a) shows minor accuracy variations across 1 to 10 prompts. While 8 prompts yield the highest accuracy (62.5%), practical test-time adaptation often precludes fine-tuning prompts on unseen data. Therefore, we select 3 prompts for our main experiments, balancing performance with practical considerations. Fig E.1(b) reveals the effect of batch size. Smaller batches (e.g., 4, 8, 16) show reduced accuracy (e.g., 60.9% at batch size 4) due to higher variance in statistics and noisy zeroth-order gradient estimates. Performance peaks at batch size 32 (62.3%), benefiting from more adaptation steps. However, increasing to 128 leads to a slight decrease (61.2%), as large batches may over-smooth feature statistics, losing domain information. For our main experiments, we select a batch size of 64 to ensure a fair comparison.
6 Conclusion
We propose Forward-Only Zeroth-Order Optimization (FOZO), a novel forward-only TTA paradigm. Compared to the latest forward-only methods, FOZO offers significant advantages: it consistently achieves higher accuracy with notably lower runtime, memory, and fewer parameters needing updates. Furthermore, both theoretical analysis and experiments demonstrate that our dynamic perturbation strategy enables faster convergence during adaptation. These advantages demonstrate that FOZO serves as a highly efficient solution, particularly suited for deployment on resource-constrained edge devices and quantized models.
Acknowledgment
This work is supported by the National Science Foundation of China under Grant 62506249, the National Major Scientific Instruments and Equipments Development Project of National Natural Science Foundation of China under Grant 8 62427820, the Natural Science Foundation of Sichuan under grant 2024NSFSC1462, and the Fundamental Research Funds for the Central Universities under grant YJ202342.